OpenStudy (anonymous):

solve for p 3(p+q)=p A.q=-2/3p B.q=-3/2p C.p=-2/3q D.p=-3/2q solve for n then find the value of b when a=3 4a=2b-7 A.-9/2 B.5/2 C.19/2 D.17 solve for r d=rt A.r=dt B.r=t/d C.d/t D.r=d-t find the width of a rectangle with a perimeter of 90 and a length of 15 A.90 B.15 C.45 D.30

OpenStudy (nincompoop):

dang, kid do you want to post your whole semester's problem and exam too?

OpenStudy (ilovecake):

lol

OpenStudy (triciaal):

solve for p 3(p+q)=p multiply each term inside by what is outside the parenthesis group like terms isolate the variable you want to one side of the equation leave everything else on the other side. do what is necessary to keep the same value . divide each side by the coefficient of the variable. 3*p +3*q = p 3p - p + 3 q = 0 2 p = -3 q p = -3/2 q do choice D solve for B then find the value of b when a=3 4a=2b-7 follow the same steps above and then substitute for a with 3 b = 4*3 + 7 divided by 2 = 19/2 choice C solve for r d=rt divide each side by t ( the coefficient of the desired variable) r = d/t choice C

OpenStudy (triciaal):

find the width of a rectangle with a perimeter of 90 and a length of 15